# SOP-Bench: Complex Industrial SOPs for Evaluating LLM Agents

### v1

@@ Appendix D Agent Prompt Templates
D.1 Function Calling Agent (FC-Agent)
@@ 7 First, carefully read and internalize the following SOP:
34 6. If at any point you’re unsure about how to proceed or if the user request doesn’t align with the SOP, state this clearly in your response and halt the process.
@@ 41 Step 1: [Description] - Outcome: [Completed/Failed/Pending]
53 Remember to adhere strictly to the SOP and provide clear, detailed explanations for each step of the process. It’s OK for the <sop_analysis> section to be quite long.’
D.2 Prompt Template for React Agent
@@ Appendix E Examples of Real SOPs
Figure 6: Examples of patient registration SOPs
@@ Appendix F SOP-Bench: Patient Intake and Registration
F.1 SOP: Patient Intake and Registration
@@ 6.2 System Updates
F.2 Sample Patient Data (Synthetically Generated)
Table 5: Sample Patient Registration Data (First Five Records)

### v2

@@ Appendix D Agent Prompt Templates
D.1. Function Calling Agent (FC-Agent)
@@ 7 First, carefully read and internalize the following SOP:
34 6. If at any point youre unsure about how to proceed or if the user request doesnt align with the SOP, state this clearly in your response and halt the process.
@@ 41 Step 1: [Description] - Outcome: [Completed/Failed/Pending]
53 Remember to adhere strictly to the SOP and provide clear, detailed explanations for each step of the process. Its OK for the <sop_analysis> section to be quite long.
D.2. Prompt Template for React Agent
@@ Appendix E Examples of Real SOPs
Figure 6. Examples of patient registration SOPs
Multi-page document excerpt showing several examples of patient registration standard operating procedures, including structured steps, headings, and form-like instructions.
@@ Appendix F SOP-Bench: Patient Intake and Registration
F.1. SOP: Patient Intake and Registration
@@ 6.2 System Updates
F.2. Sample Patient Data (Synthetically Generated)
Table 20. Sample Patient Registration Data (First Five Records)